OpenStudy (anonymous):

there is a total of 50 chickens and cows. If there are 118 legs, how many chickens are there?

OpenStudy (anonymous):

I have to use the method of substitution and linear equations.

OpenStudy (anonymous):

I CAME UP WITH: Let x be chickens let y be cows 1. 50x+y=118 2. 2x+4y=118 1. y=118-50x 2. 2x+4(118-50x)=118 is that right!?!

OpenStudy (jdoe0001):

there are 50 chickens and cows altogether, total 50 so say h = chicken c = cows h + c = 50 so we know that chicken have 2 legs, and cows have 4 legs so the amount of chicken legs are 2 times whatever "h" is 2h the amount of cows legs are 4 times whatever "c" is 4c we know there is a total of 118 legs so we could say that \(\bf 2h + 4c = 118\) notice that a cow has twice as many legs as a chicken does so one can say that \(\bf c = 2h\) so \(\bf 2h + 4c = 118\implies 2h+4(2h)=118\)

OpenStudy (jdoe0001):

hmmm ohh shoot... forgot something ahemm

OpenStudy (jdoe0001):

anyhow

OpenStudy (141823):

so cows have 4 legs, and chickens have two legs. Let cows be x and Chickens be n. 2n+4x=226 we can use the substitution property by changing n to 2(x-50) so lets enter that in the equation. 2(50-x)+4x= 118 100-x + 4x = 118 we can simplify to 2x=18 divide by 2 x=9 so if there are 9 cows, then 50 - 9 =41. Well I got 9 cows and 41 Chickens. (i could be totally wrong)

OpenStudy (jdoe0001):

I CAME UP WITH: Let x be chickens let y be cows 1. \(\bf x+y=50\) 2. 2x+4y=118

OpenStudy (141823):

well we need to substitute, since you can't solve it with two variables.

OpenStudy (141823):

so you can change the x to 2(50-y)